Write a function in satndard form that has the given four zeros.
Zeroes: 0,-2,-3,4

Answer by Theo(13342)

the equation would be:

y = (x) * (x + 2) * (x + 3) * (x - 4)

you would look for the x-values when y = 0.

the equation would becomes (x) * (x + 2) * (x + 3) * (x - 4) = 0

the standard form of this equation would be:

y = x^4 + x^3 - 14x^2 - 24x

that's the equation that would be set to 0 to get:

x^4 + x^3 - 14x^2 - 24x = 0 and factored to get:

(x) * (x + 2) * (x + 3) * (x - 4) = 0 and solved to get:

x = 0 or x = -2 or x = -3 or x = 4.

the graph looks like this: